Mariota Industries has sales of $368,520 and costs of $174,410. The company paid $31,110 in interest and $14,200 in dividends. It also increased retained earnings by $68,690 during the year. If the company's depreciation was $19,130, what was its average tax rate?

Multiple Choice

73.57%


42.39%


16.55%


34.18%


23.59%
Answer #1

Sales = $ 3,68,520

costs = ($1,74,410)

Depreciation = ($19,130)

EBIT = $1,74,980

Interest = ($31,110)

(EBT) = $1,43,870

  Tax @ 42.39 = ( $60,980 ) {Balancing figure will come here i.e This amount will arrive by below calculations.

Note: # EBT - Tax amout (?)-  Dividend = 68,690 (Earnings carried to B/S) # 1,43,870 - Tax amout (?) - 14,200 = 68,690   

# So, Tax amout = 1,43,870-68,690-14,200

= $ 60,980

  # Rate of Tax = (60,980 / 1,43,870) * 100

= 42.39 % (Rounding off to two decimals)

Dividends = ($14,200)

Earnings carried to B/S = $68,690 ( Note : Retained earnings a/c inceresed means, profit of the current year has been added to retained earnings a/c in the given problem)